40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 21 15:13:44 2017, gautam, Update, IMC, IMC input beam mode matching IMG_7220.JPGIMC_ModeMatch.pdfsingleLensSensitivity.pdfsensitivity.pdfIMCmodeMatch.m
    Reply  Tue Mar 21 21:59:48 2017, gautam, Update, IMC, IMC input beam mode matching IMG_3654.JPGIMG_3655.JPG
       Reply  Wed Mar 22 00:33:00 2017, gautam, Update, IMC, IMC length offset nulling offsetInvestigation.pdfoffset_summing_amp.pdf
          Reply  Wed Mar 22 16:58:25 2017, gautam, Update, IMC, WFS sensing matrix measurements 
             Reply  Thu Mar 23 01:44:53 2017, gautam, Update, IMC, WFS sensing matrix measurements IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfTFsummary.pdfIMC_WFS_170322.xlsx.zip
                Reply  Thu Mar 23 08:43:11 2017, rana, Update, IMC, WFS sensing matrix measurements 
                   Reply  Thu Mar 23 23:38:58 2017, gautam, Update, IMC, MC SUS damping gains stepped down 
                      Reply  Fri Mar 24 11:26:57 2017, gautam, Update, IMC, MC SUS damping gains restored 
                         Reply  Fri Mar 24 19:04:18 2017, gautam, Update, IMC, Seismic feedforward and WFS MCL_comparison.pdfseis_comparison.pdf
Message ID: 12900     Entry time: Wed Mar 22 16:58:25 2017     In reply to: 12899     Reply to this: 12901
Author: gautam 
Type: Update 
Category: IMC 
Subject: WFS sensing matrix measurements 

I've taken a bunch of transfer function measurements from the MC ASC PIT and YAW channels to the WFS error signals using the same set of DTT templates Koji used while characterizing the WFS loops a couple of months ago, before the IMC RF changes. Analysis is underway and I will post the results here shortly...

As an aside, Rana had added 10dB and 20dB gains to all of the WFS filter banks yesterday. I tried engaging the 10dB gains on the two MC2_TRANS PD loops, and this did not seem to induce any instability. I stepped both loops and saw that as expected, the 1/e times for both of these loops is about 45 seconds now (compared to ~150 seconds at the nominal gain). These have been running all day today, and the IMC seems well behaved, so I am going to leave these on for now... Jacking up the gain on the MC2_TRANS_QPD loops by 20dB induced instability - same story for the 4 WFS loops with 10dB additional gain...

ELOG V3.1.3-